# Geometric and topological rigidity of pinched submanifolds II

### v1

@@ 2.2 Isotropic curvature
for all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$. Here, we denote
[…]
where $R$ is the curvature tensor. If the strict inequality holds, we say that $(M,g)$ has positive isotropic curvature at $x$. The manifold $(M,g)$ is said to have nonnegative (or positive) isotropic curvature if it satisfies the corresponding condition at every point and for all orthonormal four-frames.
The following result is well known (see for instance [16]).
@@ 2.3 The pinching condition and the Lawson-Simons inequality
Lawson and Simons [10] proved that specific bounds on the second fundamental form for submanifolds of spheres can result to the vanishing of homology groups with integer coefficients. By employing the second variation of area, they effectively ruled out stable minimal currents in certain dimensions. Consequently, since area minimization can be performed within a homology class, this result implies the trivialization of integral homology.
[…]
@@
([6, 10, 26]) Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be an isometric immersion of a compact Riemannian manifold and $p$ be an integer such that $11\leq p\leq n-$ 1. Assume that at any point $x\in M^{n}$ and for any orthonormal basis $\{e_{1},\ldots,e_{n}\}$ of $T_{x}M$ the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$ satisfies
[…]
@@
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be an isometric immersion of a $4$-dimensional Riemannian manifold $M^{4}$ such that the inequality (italic-$\ast$1) is satisfied at a point $x\in M^{4}$ for $k=2$. Then the following assertions hold at $x$:
[…]
$\pi_{j}\circ A_{\xi}|_{V_{j}}={\langle}\xi,\eta_{j}{\rangle}Id,$
@@ 3 Proofs of the main results
We recall a key result, specifically Proposition 16 from [17], which establishes an estimate for the Bochner-Weitzenböck operator of a submanifold in terms of its second fundamental form. For four-dimensional submanifolds, this proposition can be stated as follows:
@@
If equality holds in (2) at a point $x\in M^{4}$, then the shape operator $A_{\xi}(x)$ has at most two distinct eigenvalues, with multiplicities $2$, for every unit vector $\xi\in N_{f}M(x)$.
@@
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold. Suppose that $f$ satisfies the inequality (italic-$\ast$1) at any point for $k=2$. Then, the following assertions hold:
$(i)$ The Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative, and the manifold $M^{4}$ has nonnegative isotropic curvature.
$(ii)$ Suppose that $M^{4}$ is compact and $\beta_{2}(M)\neq 0$. Then, equality holds in (italic-$\ast$1) at every point. Furthermore, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the following conditions:
@@
$\|\alpha_{13}\|^{2}+\|\alpha_{14}\|^{2}=c+{\langle}\alpha_{11},\alpha_{44}{\rangle}.$ (5)
Here, for simplicity, we denote $\alpha_{ij}=\alpha_{f}(e_{i},e_{j})$.
@@
$(ii)$ By the Hodge theorem, there exists a nontrivial harmonic 2-form $\omega\in\Omega^{2}(M)$. The Bochner-Weitzenböck formula then gives ${\langle}\mathcal{B}^{[2]}\omega,\omega{\rangle}=0$ at any point. Therefore, equality holds in inequality (2), and as a result, Proposition 7 implies that the shape operator, associated with any normal direction has at most two distinct eigenvalues, each with multiplicity $2$.
[…]
$\begin{cases}A_{\alpha}e_{1}=\rho_{\alpha}e_{1}+\kappa_{\alpha}e_{3}+\lambda_{\alpha}e_{4}\\ A_{\alpha}e_{2}=\rho_{\alpha}e_{2}+\mu_{\alpha}e_{3}+\nu_{\alpha}e_{4}\\ A_{\alpha}e_{3}=\kappa_{\alpha}e_{1}+\mu_{\alpha}e_{2}+\sigma_{\alpha}e_{3}\\ A_{\alpha}e_{4}=\lambda_{\alpha}e_{1}+\nu_{\alpha}e_{2}+\sigma_{\alpha}e_{4},\end{cases}$ (6)
@@
Since each shape operator has at most two distinct eigenvalues, each with multiplicity $2$, it follows from (6) that
@@
Using (3), (7), and the relation $\alpha_{11}+\alpha_{44}=\mathcal{H}/2$, we deduce that ($\ast$1), which now holds as equality, is equivalent to (5). Furthermore, if (4) is satisfied, the orthonormal four-frame $\{e_{i}\}_{1\leq i\leq 4}$ satisfies all the desired properties.
Suppose now that the frame $\{e_{i}\}_{1\leq i\leq 4}$ does not satisfy (4). Consider instead the orthonormal four-frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$, where $\tilde{e}_{i}=R_{\theta}e_{i},\tilde{e}_{j}=R_{\varphi}e_{j}$, for $i=1,2,j=3,4$. Here, $R_{\theta}$ and $R_{\varphi}$ denote rotations on $V_{1}=\mbox{span}\{e_{1},e_{2}\}$ and $V_{2}=\mbox{span}\{e_{3},e_{4}\}$ by angles $\theta$ and $\varphi$, respectively. For simplicity, we set $\tilde{\alpha}_{ij}=\alpha_{f}(\tilde{e}_{i},\tilde{e}_{j})$. Then, using (3), we have
@@
Since each shape operator has at most two distinct eigenvalues, each with multiplicity $2$, following a similar argument as above, we conclude that the vectors $\tilde{\alpha}_{ij}$, for $1\leq i,j\leq 4$, satisfy (3), (5) and (7). We now claim that the angles $\theta$ and $\varphi$ can be chosen such that (4) is satisfied for the frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$. Straightforward computations yield the following relations:
@@
We can select angles $\sigma_{1}$ and $\sigma_{2}$ such that the following are satisfied
@@
$0\displaystyle 2\cos\sigma_{2}{\langle}\alpha_{13}+\alpha_{24},\alpha_{14}-\alpha_{23}{\rangle}-\sin\sigma_{2}\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}-\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}=$ 0 .
By choosing $\varphi=(\sigma_{1}+\sigma_{2})/4,\theta=(\sigma_{1}-\sigma_{2})/4$, and applying (8)-(11), it follows that
@@
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold $M^{4}$. Suppose that at a point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form of $f$ at $x$ satisfies the conditions (3)-(5) in part (ii) of Proposition 8. We consider the orthonormal basis $\{\eta_{i}\}_{1\leq i\leq 6}$ of the space of 2-vectors $\Lambda^{2}T_{x}M$, satisfying $\eta_{i}\in\Lambda_{+}^{2}T_{x}M$, $\eta_{i+3}=\ast\eta_{i}$, for $1\leq i\leq 3$, and
@@
where we have set $e_{ij}=e_{i}\wedge e_{j}$. Then the following assertions hold at $x$:
@@
$\mu^{\pm}_{2}=\;$ $2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{23}\mp\alpha_{14}\|^{2},$
$\mu^{\pm}_{3}=\;$ $2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{13}\pm\alpha_{24}\|^{2},$
@@
$(ii)$ If $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at $x$, then one of the following conditions hold:
@@
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{13}=\alpha_{24}\;\;{\text{a}nd}\;\;\alpha_{14}+\alpha_{23}\neq 0,$ (ii2)
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{14}=\alpha_{23}\;\;{\text{a}nd}\;\;\alpha_{13}\neq\alpha_{24}.$ (ii3)
$(iii)$ If $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at $x$, then one of the following conditions hold:
@@
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{13}=\alpha_{24}\;\;{\text{a}nd}\;\;\alpha_{14}\neq\alpha_{23},$ (iii2)
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{14}=\alpha_{23}\;\;{\text{a}nd}\;\;\alpha_{13}+\alpha_{24}\neq 0.$ (iii3)
@@
$(i)$ By a straightforward computation using (2.1), the Gauss equation, and (3)-(5), we obtain the following
$\mathcal{B}^{[2]}(\eta_{1})=$ $\;\big{(}\|\alpha_{14}+\alpha_{23}\|^{2}+\|\alpha_{13}-\alpha_{24}\|^{2}\big{)}\eta_{1}$
@@
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{23}-\alpha_{14}\|^{2}\big{)}\eta_{2},$
@@
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}\eta_{3},$
@@
$\mathcal{B}^{[2]}(\eta_{4})=$ $\;\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}\eta_{4}$
@@
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{23}+\alpha_{14}\|^{2}\big{)}\eta_{5},$
@@
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{13}-\alpha_{24}\|^{2}\big{)}\eta_{6}.$
@@
On the other hand, by applying the Cauchy-Schwarz inequality and using (5), we obtain
@@
$=$ $\;\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}$
$+\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{23}-\alpha_{14}\|^{2}\big{)}.$
The above together with (12), implies that
$\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}=0,$
$\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{23}-\alpha_{14}\|^{2}\big{)}=0.$
This, together with (5), clearly concludes the proof of part $(ii)$. The proof of part $(iii)$ follows similarly and is therefore omitted. ∎
@@
We will need the following result.

### v2

@@ 2.2 Isotropic curvature
for all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$. Here, $R_{ijk\ell}$ are the components of the curvature tensor $R$, defined by
$R_{ijk\ell}=g\big{(}R(e_{i},e_{j})e_{k},e_{\ell}\big{)},\;1\leq i,j,k,\ell\leq 4.$
If the strict inequality holds, we say that $(M,g)$ has positive isotropic curvature at $x$. The manifold $(M,g)$ is said to have nonnegative (or positive) isotropic curvature if it satisfies the corresponding condition at every point and for all orthonormal four-frames.
The following result is well known (see, for instance, [19]).
@@ 2.3 The pinching condition and the Lawson-Simons inequality
Lawson and Simons [12] proved that specific bounds on the second fundamental form for submanifolds of spheres can result to the vanishing of homology groups with integer coefficients. By employing the second variation of area, they effectively ruled out stable minimal currents in certain dimensions. Consequently, since the existences theorems of Federer and Fleming [8] allow area minimization within a homology class, this result implies the triviality of integral homology.
The result of Lawson and Simons [12] mentioned above was later strengthened by Elworthy and Rosenberg [6, p. 71] without requiring the bound on the second fundamental form to be strict at all points of the submanifold. In this section, we state their theorem and then quote an auxiliary result from [28] that establishes the relation between our pinching condition ($\ast$1) and the inequality ($\ast\ast$5) below required in their result.
@@
([6, 12, 30]) Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be an isometric immersion of a compact Riemannian manifold. Let $p$ be an integer such that $1\leq p\leq n-1$. Assume that at any point $x\in M^{n}$ and for every orthonormal basis $\{e_{1},\ldots,e_{n}\}$ of $T_{x}M$, the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$ satisfies
$\sum_{i=1}^{p}\sum_{j=p+1}^{n}\big{(}2\|\alpha_{f}(e_{i},e_{j})\|^{2}-{\langle}\alpha_{f}(e_{i},e_{i}),\alpha_{f}(e_{j},e_{j}){\rangle}\big{)}\leq p(n-p)c.$ ($\ast\ast$)
If there exists a point at which inequality ($\ast\ast$5) is strict for every orthonormal basis of the tangent space at that point, then the homology groups satisfy
$H_{p}(M^{n};\mathbb{Z})=H_{n-p}(M^{n};\mathbb{Z})=0.$
Next, we recall the following result from [28], which establishes the relationship between the inequalities ($\ast$1) and ($\ast\ast$5) for four-dimensional submanifolds.
@@
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be an isometric immersion of a $4$-dimensional Riemannian manifold $M^{4}$ such that inequality ($\ast$1) is satisfied at a point $x\in M^{4}$ for $k=2$. Then the following assertions hold at $x$:
$(i)$ The inequality ($\ast\ast$5) holds for $p=2$ and for all orthonormal bases of $T_{x}M$. Moreover, if inequality ($\ast$1) is strict at $x$, then ($\ast\ast$5) is also strict for every orthonormal basis of $T_{x}M$.
$(ii)$ Suppose now that equality holds in ($\ast\ast$5) for some orthonormal basis $\{e_{i}\}_{1\leq i\leq 4}$ of $T_{x}M$ and $p=2$. Then there exist normal vectors $\eta_{1},\eta_{2}\in N_{f}M(x)$ such that the shape operator $A_{\xi}$ associated to any $\xi\in N_{f}M(x)$ satisfies
$\pi_{j}\circ A_{\xi}|_{V_{j}}={\langle}\xi,\eta_{j}{\rangle}Id,\;j=1,2,$
@@ 3 Proofs of the main results
We recall a key result, namely Proposition 16 from [20], which provides an estimate for the Bochner-Weitzenböck operator of a submanifold in terms of its second fundamental form. For four-dimensional submanifolds, this proposition can be stated as follows:
@@
If equality holds in (2) at a point $x\in M^{4}$, then for every unit vector $\xi\in N_{f}M(x)$, the shape operator $A_{\xi}(x)$ has at most two distinct eigenvalues, each of multiplicity $2$.
@@
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold. Suppose that $f$ satisfies inequality ($\ast$1) at every point for $k=2$. Then, the following assertions hold:
$(i)$ The Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative, and the manifold $M^{4}$ consequently has nonnegative isotropic curvature.
$(ii)$ Suppose that $M^{4}$ is compact and satisfies $\beta_{2}(M)\neq 0$. Then equality holds in ($\ast$1) at every point, and the shape operator, associated with any normal direction has at most two distinct eigenvalues, each of multiplicity $2$. Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the following conditions:
@@
$\|\alpha_{13}\|^{2}+\|\alpha_{14}\|^{2}=c+{\langle}\alpha_{11},\alpha_{44}{\rangle},$ (5)
where, for simplicity, we denote $\alpha_{ij}=\alpha_{f}(e_{i},e_{j})$.
@@
$(ii)$ By the Hodge theorem, there exists a nontrivial harmonic 2-form $\omega\in\Omega^{2}(M)$. The Bochner-Weitzenböck formula then implies
${\langle}\mathcal{B}^{[2]}\omega,\omega{\rangle}=0$
at every point. Therefore, equality holds in inequality (2), and as a result, Proposition 7 implies that the shape operator, associated with any normal direction has at most two distinct eigenvalues, each of multiplicity $2$.
On the other hand, from part $(i)$ of Lemma 6, we know that inequality ($\ast\ast$5) holds for $p=2$ at every point and for every orthonormal basis. Since $H_{2}(M;\mathbb{Z})\neq 0$, Theorem 5 implies that at each point $x\in M^{4}$ there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that equality holds in ($\ast\ast$5) for $p=2$. It follows from part $(i)$ of Lemma 6 that equality also holds in ($\ast$1) at any point. Furthermore, by choosing an orthonormal normal basis $\{\xi_{\alpha}\}_{1\leq\alpha\leq m}$ at $x\in M^{4}$ such that the mean curvature vector is $\mathcal{H}(x)=H\xi_{1}$, part $(ii)$ of Lemma 6 implies that the corresponding shape operators $A_{\xi_{\alpha}},1\leq\alpha\leq m$, are given by
$\begin{cases}A_{\xi_{\alpha}}e_{1}=\rho_{\alpha}e_{1}+\kappa_{\alpha}e_{3}+\lambda_{\alpha}e_{4},\\ A_{\xi_{\alpha}}e_{2}=\rho_{\alpha}e_{2}+\mu_{\alpha}e_{3}+\nu_{\alpha}e_{4},\\ A_{\xi_{\alpha}}e_{3}=\kappa_{\alpha}e_{1}+\mu_{\alpha}e_{2}+\sigma_{\alpha}e_{3},\\ A_{\xi_{\alpha}}e_{4}=\lambda_{\alpha}e_{1}+\nu_{\alpha}e_{2}+\sigma_{\alpha}e_{4},\end{cases}$ (6)
@@
Since each shape operator has at most two distinct eigenvalues, each of multiplicity $2$, it follows from (6) that
@@
Using (3), (7), and the relation $\alpha_{11}+\alpha_{44}=\mathcal{H}/2$, we deduce that ($\ast$1), which now holds as an equality, is equivalent to (5). Furthermore, if (4) is satisfied, the orthonormal four-frame $\{e_{i}\}_{1\leq i\leq 4}$ satisfies all the desired properties.
Suppose now that the frame $\{e_{i}\}_{1\leq i\leq 4}$ does not satisfy (4). Consider instead the orthonormal four-frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$, defined by
$\tilde{e}_{i}=R_{\theta}e_{i},\;\tilde{e}_{j}=R_{\varphi}e_{j},\;i=1,2,\;j=3,4.$
Here, $R_{\theta}$ and $R_{\varphi}$ denote rotations on $V_{1}=\mbox{span}\{e_{1},e_{2}\}$ and $V_{2}=\mbox{span}\{e_{3},e_{4}\}$ by angles $\theta$ and $\varphi$, respectively. For simplicity, we set $\tilde{\alpha}_{ij}=\alpha_{f}(\tilde{e}_{i},\tilde{e}_{j})$. Then, using (3), we have
@@
Since each shape operator has at most two distinct eigenvalues, each of multiplicity $2$, a similar argument as above shows that the vectors $\tilde{\alpha}_{ij}$, for $1\leq i,j\leq 4$, satisfy (3), (5) and (7). We now claim that the angles $\theta$ and $\varphi$ can be chosen such that (4) is satisfied for the frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$. Straightforward computations then yield the following relations:
@@
We can select angles $\sigma_{1}$ and $\sigma_{2}$ such that the following conditions are satisfied:
@@
$2\cos\sigma_{2}{\langle}\alpha_{13}+\alpha_{24},\alpha_{14}-\alpha_{23}{\rangle}-\sin\sigma_{2}\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}-\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}=0.$
By choosing
$\varphi=\frac{1}{4}(\sigma_{1}+\sigma_{2})\;\;{\text{and}}\;\;\theta=\frac{1}{4}(\sigma_{1}-\sigma_{2}),$
and applying (8)-(11), it follows that
@@
We recall that the first normal space of an isometric immersion $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ at a point $x\in M^{n}$ is defined by
$N_{1}f(x)=\operatorname{span}\left\{\alpha_{f}(X,Y):X,Y\in T_{x}M\right\}.$
@@
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold $M^{4}$. Suppose that at a point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form of $f$ at $x$ satisfies conditions (3)-(5) in part (ii) of Proposition 8. We consider the orthonormal basis $\{\eta_{i}\}_{1\leq i\leq 6}$ of the space of 2-vectors $\Lambda^{2}T_{x}M$, defined by the relations
$\eta_{i}\in\Lambda_{+}^{2}T_{x}M,\;\;\eta_{i+3}=\ast\eta_{i},\;\;\text{for}\;\;1\leq i\leq 3,$
and
@@
where $e_{ij}=e_{i}\wedge e_{j}$. Then the following assertions hold at $x$:
@@
$\mu^{\pm}_{2}=\;$ $\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+2\left(\|\alpha_{14}\|\|\alpha_{23}\|\mp{\langle}\alpha_{14},\alpha_{23}{\rangle}\right),$
$\mu^{\pm}_{3}=\;$ $\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{14}\|^{2}+2\left(\|\alpha_{13}\|\|\alpha_{24}\|\pm{\langle}\alpha_{13},\alpha_{24}{\rangle}\right),$
@@
$(ii)$ If $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at $x$, then one of the following conditions holds:
@@
$\alpha_{13}=\alpha_{24}=0,\;\alpha_{23}=\alpha_{14}\neq 0\;\;{\text{a}nd}\;\;\alpha_{44}-\alpha_{11}=2\rho\,\alpha_{14},\;\rho\in\mathbb{R},$ (ii2)
$\alpha_{14}=\alpha_{23}=0,\;\alpha_{24}=-\alpha_{13}\neq 0\;\;{\text{a}nd}\;\;\alpha_{44}-\alpha_{11}=2\rho\,\alpha_{13},\;\rho\in\mathbb{R}.$ (ii3)
$(iii)$ If $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at $x$, then one of the following conditions holds:
@@
$\alpha_{13}=\alpha_{24}=0,\;\alpha_{23}=-\alpha_{14}\neq 0\;\;{\text{a}nd}\;\;a_{44}-\alpha_{11}=2\rho\,\alpha_{14},\;\rho\in\mathbb{R},$ (iii2)
$\alpha_{14}=\alpha_{23}=0,\;\alpha_{24}=\alpha_{23}\neq 0\;\;{\text{a}nd}\;\;\alpha_{44}-\alpha_{11}=2\rho\,\alpha_{13},\;\rho\in\mathbb{R}.$ (iii3)
In addition, in cases (ii2), (ii3), (iii2), and (iii3), the immersion $f$ has flat normal bundle at $x$, and its first normal space satisfies
$1\leq\dim N_{1}f(x)\leq 2.$
@@
$(i)$ By a straightforward computation using (2.1), the Gauss equation, and (3)-(5), we obtain the following:
$\mathcal{B}^{[2]}(\eta_{1})=$  $\;2\big{(}2c+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-\|\alpha_{13}\|^{2}-\|\alpha_{14}\|^{2}+{\langle}\alpha_{14},\alpha_{23}{\rangle}-{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{1}$
@@
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{14}\|^{2}-2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{2}$
$=$ $\;a_{1}^{+}\eta_{1}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|\|\alpha_{23}\|-2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{2},$
@@
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{13}\|^{2}+2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{3}$
$=$ $\;a_{2}^{+}\eta_{1}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{14}\|^{2}+2\|\alpha_{13}\|\|\alpha_{24}\|+2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{3},$
@@
$\mathcal{B}^{[2]}(\eta_{4})=$  $\;2\big{(}2c+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-\|\alpha_{13}\|^{2}-\|\alpha_{14}\|^{2}-{\langle}\alpha_{14},\alpha_{23}{\rangle}+{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{4}$
@@
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{14}\|^{2}+2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{5}$
$=$ $\;a_{1}^{-}\eta_{4}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|\|\alpha_{23}\|+2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{5},$
@@
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{13}\|^{2}-2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{6}$
$=$ $\;a_{2}^{-}\eta_{4}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{14}\|^{2}+2\|\alpha_{13}\|\|\alpha_{24}\|-2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{6}.$
@@
On the other hand, by applying the Cauchy-Schwarz inequality, we obtain
@@
$=$  $\;2\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{14}\|^{2}+\|\alpha_{13}\|\|\alpha_{24}\|+{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}$
$+2\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+\|\alpha_{14}\|\|\alpha_{23}\|-{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}.$
The above, together with (12), implies that
$\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{14}\|^{2}+\|\alpha_{13}\|\|\alpha_{24}\|+{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}=0,$
$\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+\|\alpha_{14}\|\|\alpha_{23}\|-{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}=0.$
This clearly concludes the proof of part $(ii)$. The proof of part $(iii)$ proceeds similarly and is therefore omitted.
The Ricci equation states that the normal curvature tensor $R^{\perp}$ of the submanifold $f$ is given by
$R^{\perp}(X,Y)\xi=\alpha_{f}(X,A_{\xi}Y)-\alpha_{f}(A_{\xi}X,Y)$
for all $X,Y\in\mathcal{X}(M)$ and $\xi\in\Gamma(N_{f}M)$. We claim that
$\alpha_{f}(e_{i},A_{\xi}e_{j})-\alpha_{f}(A_{\xi}e_{i},e_{j})=0\;\;\text{for all}\;\;1\leq i,j\leq 4,$ (13)
which clearly implies that the immersionff has flat normal bundle at the point $x$.
We prove (13) in the case where (ii2) holds, since the same argument applies to the other cases. Let $(h_{ij})$ denote the matrix of the shape operator $A_{\xi}$ with respect to the four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$. Then it follows from part (ii2) of Lemma 9 that
$h_{13}=h_{24}=0,\;h_{23}=-h_{14}\neq 0\;\;\text{and}\;\;h_{44}-h_{11}=2\rho\,h_{14}.$
That (13) holds follows directly from these relations together with part (ii2) of Lemma 9.
Finally, the inequality on the dimension of the first normal space at $x$ follows directly from parts (ii2), (ii3), (iii2) and (iii3) of Lemma 9. ∎
@@
The following result will be used in the sequel.